OpenStudy (anonymous):

A wallet containing £40 has three times as many $1 notes as 5 notes. Find the number of each kind?

OpenStudy (hoblos):

the first is in £ and the others are in $ do you know how to convert??

OpenStudy (anonymous):

no that is a typo. the currency is all the same.

OpenStudy (cwrw238):

oh - thats why i got confused!!

OpenStudy (hoblos):

0k let x=number of $1 notes y=number of $5 notes x+5y =40 x=3y -------> replace this in the first equation 3y+5y=40 y=5 x=3y=15

OpenStudy (hoblos):

number of $1 notes is 15 number of $5 notes is 5

OpenStudy (anonymous):

how did you come up with those systems?

OpenStudy (hoblos):

"three times as many $1 notes as 5 notes" means s=3y this is obvious "wallet containing $40" means that the sum of all the money is $40 so x*($1) + y*($5) =$40

OpenStudy (anonymous):

are you a genius???

OpenStudy (hoblos):

no not really :P

OpenStudy (anonymous):

one last thing, I thught because there were 3 times more £1 notes "x" then £5 notes "y" that 3x=y.

OpenStudy (hoblos):

what i understand from this is that x=3y !! well you better ask somebody else about to make sure

OpenStudy (anonymous):

but how does that show that there are 3 times more x's then y's?